# A mesoscopic model for the rheology of dilute and semidilute solutions of wormlike micelles

### v1

@@ Abstract
The concept of a ‘persistent worm’ is introduced, representing the smallest possible length of a wormlike micelle, and modelled by a bead-spring chain with sticky beads at the ends. Persistent worms are allowed to combine with each other at their sticky ends to form wormlike micelles with a distribution of lengths, and the semiflexibility of a wormlike micelle is captured with a bending potential between springs, both within and across persistent worms that stick to each other. Multi-particle Brownian dynamics simulations of such polydisperse and ‘polyflexible’ wormlike micelles, with hydrodynamic interactions included and coupled with reversible scission/fusion of persistent worms, are used to investigate the static and dynamic properties of wormlike micellar solutions in the dilute and unentangled semidilute concentration regimes. The influence of the sticker energy and persistent worm concentration are examined and simulations are shown to validate theoretical mean-field predictions of the universal scaling with concentration of the chain length distribution of linear wormlike micelles, independent of the sticker energy. The presence of wormlike micelles that form rings is shown not to affect the static properties of linear wormlike micelles, and mean-field predictions of ring length distributions are validated. Linear viscoelastic storage and loss moduli are computed and the unique features in the intermediate frequency regime compared to those of homopolymer solutions are highlighted. The distinction between Rouse and Zimm dynamics in wormlike micelle solutions is elucidated, with a clear identification of the onset of the screening of hydrodynamic interactions with increasing concentration.
@@ I Introduction
Between the two extremes of continuum models and coarse-grained mesoscopic models for wormlike micelle solutions lie phenomenological models that are ‘microstructurally’ inspired, and which aim to predict both the shapes of flow curves and the existence of banded structures in shear flow [69, 71, 77, 1, 25, 26]. In contrast to vorticity banding, which is typically observed in shear flows of dilute wormlike micellar solutions and requires the existence of a reentrant flow curve, shear banding is observed in semidilute entangled and concentrated solutions, and requires the shear rate to be a multi-valued function of shear stress over a range of shear stresses [46, 17]. Phenomenological models for semidilute entangled and concentrated solutions, such as the VCM model [71, 77] and Brownian dynamics simulations [54] of refined versions of these models [1] are able to capture the non-monotonic dependence of shear stress on shear rate. These studies, which are built on the representation of micelles by two species of dumbbells, short and long, where two short dumbbells can combine to form a long one and the long dumbbell can break into two short ones, suggest that the high shear rate band consists of the short micelle species while the low shear rate band consists of the longer micelle species. This is consistent with experimental observations of differences in birefringence in the two banded regions in Couette flow that are at different shear rates but at the same shear stress [16]. Though examination of the semidilute entangled and concentrated regimes is outside the scope of this work, it is worth pointing out that this insight into the molecular composition of the shear bands obtained with models that are highly coarse-grained, is already quite valuable, and suggests that much greater understanding could be gained by the development of more fine-grained models which are capable of capturing wormlike micellar structure and dynamics on the molecular scale more accurately and that can predict the occurrence of banding. Such models are, however, currently lacking.
@@
In the case of dilute wormlike micellar solutions, in two recent groundbreaking papers [25, 26], Hommel and Graham have introduced the reactive rod model, which belongs to the class of phenomenological models that are strongly motivated by physics on the microscopic scale, and which is capable of capturing several of the key experimentally observed rheological signatures in these solutions, such as shear thickening and thinning in simple shear flow [25], and vorticity banding and the presence of finger-like instabilities in circular Couette flow [26]. In this model, wormlike micelles are represented as rods that can combine together by fusion into longer rods, which in turn can breakdown into shorter rods by scission. Both the fusion and scission can happen spontaneously or be induced by flow. By solving coupled evolution equations for the average orientation of rods and the collective length of the micelles, the model computes the micelle contribution to solution stress, which is then combined with the equations for the conservation of mass and momentum to solve for macroscopic observables in any geometry. These papers represent a significant advance in the description of dilute wormlike micellar solutions and the reactive rod model will undoubtedly provide greater insight into experimental observations in other flow fields. Nevertheless, there are some aspects of wormlike micelle structure and dynamics that the model does not take into account.
@@
Given the vast range of length and time scales involved, carrying out simulations at the level of individual surfactant molecules in order to model the rheological behaviour of wormlike micelle solutions is still not computationally feasible, and scales that are relevant to rheological properties can only be probed meaningfully through coarse-grained simulations at the mesoscopic level [48, 49]. There have been many mesoscopic models developed to describe the rheological behaviour of wormlike micellar solutions based on a variety of different numerical methods ranging from non-equilibrium molecular dynamics [35, 8, 34, 47] to Brownian dynamics simulations [33, 48, 30, 49, 1, 54], and predictions of a number of linear and nonlinear rheological properties in shear flow have been made. All these approaches have as their key component the fusion of individual ‘indivisible’ units that represent the shortest possible length of a wormlike micelle into longer linear or ring-like wormlike micelles, which subsequently breakup by scission, possibly all the way back down to the individual indivisible units. These indivisible units are typically beads [35, 8, 33, 47, 30], dumbbells [1, 54] or rods [48, 49] and a range of different potentials and algorithms have been used to determine the processes of fusion and scission. In nearly all these cases, the size of the single indivisible unit is equal to one persistence length of the wormlike micelle. The only exceptions are the FENE-CB model of Kröger [33, 34] and rod model of Padding et. al. [48, 49], where a bending potential is used between beads in the former case and rods in the latter, to account for semiflexibility. In neither of these cases, however, has the effect of the bending rigidity of wormlike micelles on their rheological behaviour been studied systematically. Models that assume that the length of the indivisible unit is equal to a persistence length do not account for semiflexibility since the wormlike micelles formed by their fusion are intrinsically flexible in nature. Additionally, nearly all existing mesoscopic models neglect the presence of hydrodynamic interactions. Early non-equilibrium molecular dynamics simulations did account for the presence of hydrodynamic interactions by including solvent molecules explicitly [35, 8]. However, in order to examine the influence of hydrodynamic interactions, it is necessary to compare results of simulations with and without their presence, which has not been done so far. While neglecting hydrodynamic interactions is well justified for solutions that are sufficiently concentrated such that they are screened [59], there has not been a meticulous investigation so far to determine the critical concentration beyond which hydrodynamic interactions may be neglected. In the context of the accurate prediction of the rheological behaviour of dilute and semidilute homopolymer solutions, it is now well established that it is vitally important to include hydrodynamic interactions [53]. The aim of the present work is to introduce a mesoscopic model that is valid for finite concentrations, which can account for both semiflexibility and hydrodynamic interactions, and as a consequence, distinguish between Rouse and Zimm dynamics in wormlike micelle solutions and enable the determination of the critical concentration beyond which hydrodynamic interactions are screened.
@@
The shortest possible length of a wormlike micelle is clearly determined by entropic and enthalpic considerations, and would depend on the chemistry and geometry of the surfactant molecules, the chemistry of the solvent molecules and ions present in solution, and on their mutual interactions. It seems reasonable to argue that the resultant shortest possible length need not necessarily be equal to exactly one persistent length. In this work, we introduce the concept of a ‘persistent worm’, which is the shortest possible length of a wormlike micelle, that can in principle be any fraction of the actual persistent length of the wormlike micelle. The persistent worm is modelled as a bead-spring chain with an arbitrary number of beads, and the terminal beads are made ‘sticky’, i.e., they can associate with other sticky beads to form pairs of stuck beads and thus form a wormlike micelle. A bending potential, with which the persistent length can be tuned, is imposed between the springs within a persistent worm, and when two persistent worms combine, the same bending potential is imposed across the two persistent worms in order to maintain a uniform degree of bending stiffness along the entire backbone of a wormlike micelle formed by the fusion of persistent worms. Since the degree of flexibility of a polymer chain is determined by the ratio of the persistent length to the contour length, this implies that the current model leads to a polydisperse and ‘polyflexible’ solution of wormlike micelles. Hydrodynamic interactions have been included within the framework of HOOMD-Blue, using an efficient algorithm for their fast computation [23]. With this mesoscopic model, Brownian dynamics simulation predictions of equilibrium properties and the linear viscoelasticity of wormlike micelle solutions, across a range of concentrations from the dilute to the unentangled semidilute regime, are made as a function of model parameters. The model is validated by comparison of static property predictions with the classical results of mean-field and scaling theories, and the influence of hydrodynamic interactions on storage and loss moduli is explored for the first time.
@@ II.2.1 Lengths, concentrations and distributions
[…]
@@
$\psi_{\text{lin}}(L)\Delta L$ $=\frac{n^{\text{lin}}_{\text{L}+\Delta\text{L}}-n^{\text{lin}}_{\text{L}}}{\mathcal{N}^{\text{lin}}_{\text{wlm}}}$
$\text{and},\,\,\psi_{\text{R}}(L)\Delta L$ $=\frac{n^{\text{R}}_{\text{L}+\Delta\text{L}}-n^{\text{R}}_{\text{L}}}{\mathcal{N}^{\text{R}}_{\text{wlm}}}$ (10)
@@ II.2.2 The pairwise bead-bead interaction potential
$\frac{U_{\mu\nu}^{\text{SDK}}}{k_{\text{B}}T}=\left\{\begin{array}[]{l l}4\!\left[\!\left(\dfrac{\sigma}{r_{\mu\nu}}\right)^{12}-\left(\dfrac{\sigma}{r_{\mu\nu}}\right)^{6}+\dfrac{1}{4}\right]-\epsilon\,;&r_{\mu\nu}\leq 2^{1/6}\sigma\\[15.0pt] \dfrac{1}{2}\,\epsilon\!\left[\cos\!\left(\!\alpha\left(\dfrac{r_{\mu\nu}}{\sigma}\right)^{2}+\beta\!\right)-1\right];&\!\!2^{1/6}\sigma\leq r_{\mu\nu}\leq r_{\rm c}\\[15.0pt] 0&r_{\mu\nu}\geq r_{\rm c}\end{array}\right.$ (16)
@@
The repulsive part of the SDK potential is identical to a truncated Lennard-Jones potential, while the attractive part is represented by a cosine function. The minimum of the attractive well-depth occurs at a distance of $r_{\mu\nu}=2^{1/6}\sigma$ and $\epsilon$ is the well-depth of the potential in units of $k_{B}T$. In terms of the length scale $l_{H}=\sqrt{k_{\text{B}}T/H}$ (used in this study for non-dimensionalising lengths), $\sigma/l_{H}$ is set equal to 1 in all the simulations reported here. Following the arguments laid out in the recent work by Santra et al. [63], the cutoff radius $r_{c}$ is set equal to $1.82\,\sigma$. The constants $\alpha$ and $\beta$ are determined by the boundary conditions: $\text{U}_{\mu\nu}^{\text{SDK}}=0$ at $r_{\mu\nu}=r_{c}$ and $\text{U}_{\mu\nu}^{\text{SDK}}=-\epsilon$ at $r_{\mu\nu}=2^{1/6}\sigma$. Using these conditions, the values of $\alpha$ and $\beta$ are computed to be 1.530633312 and 1.213115524, respectively. Amongst the many advantages of the SDK potential, is its ability to interpolate between good solvent and poor solvent conditions by varying the well depth $\epsilon$ [63]. Here, the well-depth corresponding to interactions between backbone-backbone and backbone-sticker bead pairs is denoted by $\epsilon_{bb}$, and both are considered to be purely repulsive, i.e., $\epsilon_{bb}=0$, corresponding to the good solvent limit. The well-depth corresponding to associative interactions between sticker-sticker beads is denoted by $\epsilon_{st}$, and is typically chosen to be strongly attractive, with $\epsilon_{st}\geq 1$.
@@
The choice of using the SDK potential here to describe associative interactions implies that while there is a free energy barrier to the scission of two sticky beads that are stuck together, there is no barrier to their fusion if they are within cut-off distance of each other. Several previous studies have included a free energy barrier to fusion, arguing that the recombination of chains maybe a complex activated process [29, 30, 47, 49]. For simplicity, in this preliminary study that aims to introduce a new mesoscopic model for wormlike micelles, an activation barrier for fusion has been neglected. Depending on the performance of the model in its ability to capture rheological signatures of wormlike micelle solutions, an activation barrier can be introduced subsequently if this is considered necessary to improve its predictions. The occurrence of a scission/fusion event is determined by a Monte Carlo scheme, which is described in greater detail below.

### v2

@@ Abstract
The concept of a ‘persistent worm’ is introduced, representing the smallest possible length of a wormlike micelle, and modelled by a bead-spring chain with sticky beads at the ends. Persistent worms are allowed to combine with each other at their sticky ends to form wormlike micelles with a distribution of lengths, and the semiflexibility of a wormlike micelle is captured with a bending potential between springs, both within and across persistent worms that stick to each other. Multi-particle Brownian dynamics simulations of such polydisperse and ‘polyflexible’ wormlike micelles, with hydrodynamic interactions included and coupled with reversible scission/fusion of persistent worms, are used to investigate the static and dynamic properties of wormlike micellar solutions in the dilute and unentangled semidilute concentration regimes. The influence of the sticker energy and persistent worm concentration are examined and simulations are shown to validate theoretical mean-field predictions of the universal scaling with concentration of the chain length distribution of linear wormlike micelles, independent of the sticker energy. The presence of wormlike micelles that form rings is shown not to affect the static properties of linear wormlike micelles, and mean-field predictions of ring length distributions are validated. Linear viscoelastic storage and loss moduli are computed and the unique features in the intermediate frequency regime compared to those of homopolymer solutions are highlighted. The inclusion of hydrodynamic interactions enables the distinction between Rouse and Zimm dynamics in wormlike micelle solutions to be elucidated. While the neglect of hydrodynamic interactions is justified in the case of entangled wormlike micelle solutions, here the scaled concentration $c/c^{*}$ at which the onset of the screening of hydrodynamic interactions occurs with increasing concentration, as a function of the scission energy, is clearly identified.
@@ I Introduction
Between the two extremes of continuum models and coarse-grained mesoscopic models for wormlike micelle solutions lie phenomenological models that are ‘microstructurally’ inspired, and which aim to predict both the shapes of flow curves and the existence of banded structures in shear flow [72, 74, 80, 1, 26, 27]. In contrast to vorticity banding, which is typically observed in shear flows of dilute wormlike micellar solutions and requires the existence of a reentrant flow curve, shear banding is observed in semidilute entangled and concentrated solutions, and requires the shear rate to be a multi-valued function of shear stress over a range of shear stresses [49, 17].
@@
The non-monotonic dependence of shear stress on shear rate is captured by phenomenological models for semidilute entangled and concentrated solutions, such as the VCM model [74, 80] (where entanglement effects are treated implicitly through an empirically tuneable parameter) and Brownian dynamics simulations [57] of refined versions of these models [1]. These studies, which are built on the representation of micelles by two species of dumbbells, short and long, where two short dumbbells can combine to form a long one and the long dumbbell can break into two short ones, suggest that the high shear rate band consists of the short micelle species while the low shear rate band consists of the longer micelle species. This is consistent with experimental observations of differences in birefringence in the two banded regions in Couette flow that are at different shear rates but at the same shear stress [16]. Though examination of the semidilute entangled and concentrated regimes is outside the scope of this work, it is worth pointing out that this insight into the molecular composition of the shear bands obtained with models that are highly coarse-grained, is already quite valuable, and suggests that much greater understanding could be gained by the development of more fine-grained models which are capable of capturing wormlike micellar structure and dynamics on the molecular scale more accurately and that can predict the occurrence of banding. Such models are, however, currently lacking.
@@
In the case of dilute wormlike micellar solutions, the reactive rod model was originally introduced by Dutta and Graham [20] and further refined to better capture experimental results in two recent groundbreaking papers by Hommel and Graham [26, 27]. This model belongs to the class of phenomenological models that are strongly motivated by physics on the microscopic scale, and which is capable of capturing several of the key experimentally observed rheological signatures in these solutions, such as shear thickening and thinning in simple shear flow [26], and vorticity banding and the presence of finger-like instabilities in circular Couette flow [27]. In this model, wormlike micelles are represented as rods that can combine together by fusion into longer rods, which in turn can breakdown into shorter rods by scission. Both the fusion and scission can happen spontaneously or be induced by flow. By solving coupled evolution equations for the average orientation of rods and the collective length of the micelles, the model computes the micelle contribution to solution stress, which is then combined with the equations for the conservation of mass and momentum to solve for macroscopic observables in any geometry. These papers represent a significant advance in the description of dilute wormlike micellar solutions and the reactive rod model will undoubtedly provide greater insight into experimental observations in other flow fields. Nevertheless, there are some aspects of wormlike micelle structure and dynamics that the model does not take into account.
@@
Given the vast range of length and time scales involved, carrying out simulations at the level of individual surfactant molecules in order to model the rheological behaviour of wormlike micelle solutions is still not computationally feasible, and scales that are relevant to rheological properties can only be probed meaningfully through coarse-grained simulations at the mesoscopic level [51, 52]. There have been many mesoscopic models developed to describe the rheological behaviour of wormlike micellar solutions based on a variety of different numerical methods ranging from non-equilibrium molecular dynamics [36, 8, 35, 50] to Brownian dynamics simulations [34, 51, 31, 52, 1, 57], and predictions of a number of linear and nonlinear rheological properties in shear flow have been made. All these approaches have as their key component the fusion of individual ‘indivisible’ units that represent the shortest possible length of a wormlike micelle into longer linear or ring-like wormlike micelles, which subsequently breakup by scission, possibly all the way back down to the individual indivisible units. These indivisible units are typically beads [36, 8, 34, 50, 31], dumbbells [1, 57] or rods [51, 52] and a range of different potentials and algorithms have been used to determine the processes of fusion and scission. In nearly all these cases, the size of the single indivisible unit is equal to one persistence length of the wormlike micelle. The only exceptions are the FENE-CB model of Kröger [34, 35] and rod model of Padding et. al. [51, 52], where a bending potential is used between beads in the former case and rods in the latter, to account for semiflexibility. In neither of these cases, however, has the effect of the bending rigidity of wormlike micelles on their rheological behaviour been studied systematically. Models that assume that the length of the indivisible unit is equal to a persistence length do not account for semiflexibility since the wormlike micelles formed by their fusion are intrinsically flexible in nature.
@@
The single indivisible unit of a wormlike micelle is clearly determined by entropic and enthalpic considerations, and would depend on the chemistry and geometry of the surfactant molecules, the chemistry of the solvent molecules and ions present in solution, and on their mutual interactions. [32, 46] It seems reasonable to argue that the resultant shortest possible length need not necessarily be equal to exactly one persistent length. In contrast to the previous work, we introduce the concept of a ‘persistent worm’, which is the shortest possible length of a wormlike micelle, that can in principle be any fraction of the actual persistent length of the wormlike micelle. The persistent worm is modelled as a bead-spring chain with an arbitrary number of beads, and the terminal beads are made ‘sticky’, i.e., they can associate with other sticky beads to form pairs of stuck beads and thus form a wormlike micelle. A bending potential, with which the persistent length can be tuned, is imposed between the springs within a persistent worm, and when two persistent worms combine, the same bending potential is imposed across the two persistent worms in order to maintain a uniform degree of bending stiffness along the entire backbone of a wormlike micelle formed by the fusion of persistent worms. Since the degree of flexibility of a polymer chain is determined by the ratio of the persistent length to the contour length, this implies that the current model leads to a polydisperse and ‘polyflexible’ solution of wormlike micelles.
Additionally, nearly all existing mesoscopic models neglect the presence of hydrodynamic interactions. Early non-equilibrium molecular dynamics simulations did account for the presence of hydrodynamic interactions by including solvent molecules explicitly [36, 8]. However, in order to examine the influence of hydrodynamic interactions, it is necessary to compare results of simulations with and without their presence, which has not been done so far. While neglecting hydrodynamic interactions is well justified for solutions that are sufficiently concentrated such that they are screened [62], there has not been a meticulous investigation so far to determine the critical concentration beyond which hydrodynamic interactions may be neglected. In the context of the accurate prediction of the rheological behaviour of dilute and semidilute homopolymer solutions, it is now well established that it is vitally important to include hydrodynamic interactions [56].
In the present work, a mesoscopic model is introduced that is valid for finite concentrations, which can account for both semiflexibility and hydrodynamic interactions, and as a consequence, distinguish between Rouse and Zimm dynamics in wormlike micelle solutions and enable the determination of the critical concentration beyond which hydrodynamic interactions are screened. Essentially, hydrodynamic interactions have been included within the framework of HOOMD-Blue, using an efficient algorithm for their fast computation [24].
With this mesoscopic model, Brownian dynamics simulation predictions of equilibrium properties and the linear viscoelasticity of wormlike micelle solutions, across a range of concentrations from the dilute to the unentangled semidilute regime, are made as a function of model parameters. The model is validated by comparison of static property predictions with the classical results of mean-field and scaling theories, and the influence of hydrodynamic interactions on storage and loss moduli is explored for the first time.
@@ II.2.1 Lengths, concentrations and distributions
$\text{and},\quad$ $\mathcal{N}_{\text{wlm}}=\mathcal{N}^{\text{lin}}_{\text{wlm}}+\mathcal{N}^{\text{R}}_{\text{wlm}}$ (5)
@@
$\psi_{\text{lin}}(L)\Delta L$ $=\frac{n^{\text{lin}}_{L+\Delta L}-n^{\text{lin}}_{L}}{\mathcal{N}^{\text{lin}}_{\text{wlm}}},$
$\text{and}\quad\psi_{\text{R}}(L)\Delta L$ $=\frac{n^{\text{R}}_{L+\Delta L}-n^{\text{R}}_{L}}{\mathcal{N}^{\text{R}}_{\text{wlm}}}$ (10)
@@ II.2.2 The pairwise bead-bead interaction potential
$\frac{U_{\mu\nu}^{\text{SDK}}}{k_{\text{B}}T}=\left\{\begin{array}[]{l l}\hskip-2.84526pt4\left[\left(\dfrac{\sigma}{r_{\mu\nu}}\right)^{12}-\left(\dfrac{\sigma}{r_{\mu\nu}}\right)^{6}+\dfrac{1}{4}\right]-\epsilon\,;&r_{\mu\nu}\leq 2^{1/6}\sigma\\[15.0pt] \hskip-2.84526pt\dfrac{1}{2}\,\epsilon\left[\cos\left(\alpha\left(\dfrac{r_{\mu\nu}}{\sigma}\right)^{2}+\beta\right)-1\right];&\hskip-14.22636pt2^{1/6}\sigma\leq r_{\mu\nu}\leq r_{\rm c}\\[15.0pt] \hskip-2.84526pt0.&r_{\mu\nu}\geq r_{\rm c}\end{array}\right$. (16)
@@
The repulsive part of the SDK potential is identical to a truncated Lennard-Jones potential, while the attractive part is represented by a cosine function. The minimum of the attractive well-depth occurs at a distance of $r_{\mu\nu}=2^{1/6}\sigma$ and $\epsilon$ is the well-depth of the potential in units of $k_{B}T$. In terms of the length scale $l_{H}=\sqrt{k_{\text{B}}T/H}$ (used in this study for non-dimensionalising lengths), $\sigma/l_{H}$ is set equal to 1 in all the simulations reported here. Following the arguments laid out in the recent work by Santra et al. [66], the cutoff radius $r_{c}$ is set equal to $1.82\,\sigma$. The constants $\alpha$ and $\beta$ are determined by the boundary conditions: $\text{U}_{\mu\nu}^{\text{SDK}}=0$ at $r_{\mu\nu}=r_{c}$ and $\text{U}_{\mu\nu}^{\text{SDK}}=-\epsilon$ at $r_{\mu\nu}=2^{1/6}\sigma$. Using these conditions, the values of $\alpha$ and $\beta$ are computed to be 1.530633312 and 1.213115524, respectively.
Figure 2: The Soddemann-Dünweg-Kremer (SDK) potential ($U_{\mu\nu}^{\mathrm{SDK}}/k_{\text{B}}T$), defined in Eq. (16), plotted as a function of the radial distance $r_{\mu\nu}$ (in units of $l_{H}$), for different well depths $\epsilon_{st}$.
Amongst the many advantages of the SDK potential is its ability to interpolate between good solvent and poor solvent conditions by varying the well depth $\epsilon$ [66]. Here, the well-depth corresponding to interactions between backbone-backbone and backbone-sticker bead pairs is denoted by $\epsilon_{bb}$, and both are considered to be purely repulsive, i.e., $\epsilon_{bb}=0$, corresponding to the good solvent limit. The well-depth corresponding to associative interactions between sticker-sticker beads is denoted by $\epsilon_{st}$, and is typically chosen to be strongly attractive, with $\epsilon_{st}\geq 1$. The shape of the SDK potential for various values of the well depth $\epsilon_{st}$ is illustrated in Fig. 2.
In reality, the distribution of wormlike micelle lengths is governed by a balance between entropic and enthalpic factors, with the equilibrium micelle length arising from a trade-off between the enthalpic cost of forming hemispherical endcaps and the entropic penalty of maintaining long micelles. Forming an endcap requires a distortion in the packing of surfactant molecules and a loss of favourable tail–tail interactions, which introduces an energy penalty, typically termed the endcap energy [46, 43, 32]. In the present model, where the complex energetics of scission and fusion is captured with a phenomenological SDK interaction potential between the “sticky” ends of persistent worms, the well depth $\epsilon_{st}$ acts as a surrogate for the endcap energy. It quantifies the energy lost when two sticky ends fuse to form a longer chain, thereby eliminating two endcaps, and conversely, the creation of two new endcaps through scission leads to an increase in the micelle energy by $\epsilon_{st}$. This approach is consistent with previous coarse-grained models, where similar short-range potentials have been used to capture complex fusion–scission dynamics in micellar systems [10, 75, 76, 51, 30, 52]. Within this coarse-grained framework, as discussed in greater detail in  Sec. III.3.1, using arguments based on minimizing the Helmholtz free energy, mean-field theories predict the distribution of chain lengths and the mean micelle length’s dependence on concentration and scission energy. As demonstrated subsequently in  Sec. III.3.2, the use of the SDK potential effectively captures the thermodynamic cost of endcap formation and removal, and recovers the theoretical predictions of equilibrium static properties based on free energy minimization arguments.
@@
Further, the choice of using the SDK potential to describe associative interactions implies that while there is a free energy barrier to the scission of two sticky beads that are stuck together, there is no barrier to their fusion if they are within cut-off distance of each other. Several previous studies have included a free energy barrier to fusion, arguing that the recombination of chains maybe a complex activated process [30, 31, 50, 52]. For simplicity, in this preliminary study that aims to introduce a new mesoscopic model for wormlike micelles, an activation barrier for fusion has been neglected. Depending on the performance of the model in its ability to capture rheological signatures of wormlike micelle solutions, an activation barrier can be introduced subsequently if this is considered necessary to improve its predictions. The occurrence of a scission/fusion event is determined by a Monte Carlo scheme, which is described in greater detail below.